# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ Appendix E Other proofs
Since $j^{m}\in J_{\neg Q}$, the agent has no information about $j^{m}$ other than that it is in $J_{\neg Q}$. This means that for all $j\in J_{\neg Q}$ and $t\in V_{j}$, $j^{m}$ is conditionally (under the condition of $j^{m}\in J_{\neg Q}$) independent of $y_{t}$. We proceed by case analysis.
Case 1: For all $j\in J_{\neg Q},t\in V_{j}$, we have $y_{t}=0$. In particular, this holds for $j=j^{m}$, and we know there exists at least one $t\in V_{j^{m}}$ since $|V_{j^{m}}|>0$. Then $y_{t}\neq\pi^{m}(x_{t})$, so $\mu(x_{t},y_{t})=0$ and thus $\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{and }|V_{j^{m}}|>0\right]=1$.
Case 2: There exists $j\in J_{\neg Q},t\in V_{j}$ with $y_{t}=1$. Then $\mu(x_{t},y_{t})=0$ unless $j=j^{m}$, so
$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\geq$ $\ \Pr\Big{[}\mu(x_{t},y_{t})=0\mid j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
$=$ $\ \Pr\Big{[}j\neq j^{m}\mid j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
Conditioned on $j^{m}\in J_{\neg Q}$, $j^{m}$ is uniformly distributed across $J_{\neg Q}$, so
$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\geq 1-\frac{1}{|J_{\neg Q}|}\geq 1-\frac{1}{f(T)-g(T)}$
@@
$\quad\quad\quad=\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\cdot\Pr\Big{[}j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
$\quad\quad\quad\geq\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
@@
$\lim_{T\to\infty}\alpha(T)=$ $\ \lim_{T\to\infty}\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
$=$ $\ (1-0)(1-0-0)$
[…]
Part 4: Putting it all together. Since $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq 1$ always, we have
$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]=$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]$
$+$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})\neq 0\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\neq 0\right]$
$\leq 0$ $\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]+1\cdot\left(1-\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\right)$
$\leq 1$ $-\alpha(T)$
@@
$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
$=$ $\ 1-\operatorname*{\mathbb{E}}_{j^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]$
$\geq$ $\ \alpha(T)$
Therefore $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]\geq\lim_{T\to\infty}\alpha(T)=1$, as required. ∎

### v5

@@ Appendix E Properties of local generalization
Case 1: For all $j\in J_{\neg Q}$ and $t\in V_{j}$, we have $y_{t}=0$. In particular, this holds for $j=j^{m}$, and we know there exists at least one $t\in V_{j^{m}}$ since $|V_{j^{m}}|>0$. Then $y_{t}\neq\pi^{m}(x_{t})$, so $\mu(x_{t},y_{t})=0$ and thus $\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{and }|V_{j^{m}}|>0\text{ and }y_{t}=0\ \forall j\in J_{\neg Q},t\in V_{j}\right]=1$.
Case 2: There exists $j\in J_{\neg Q}$ and $t\in V_{j}$ with $y_{t}=1$. Then $\mu(x_{t},y_{t})=0$ unless $j=j^{m}$, so
$\ \Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\right]$
$\geq$ $\ \Pr\Big{[}\mu(x_{t},y_{t})=0\ \Big{|}\ j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\Big{]}$
$=$ $\ \Pr\Big{[}j\neq j^{m}\ \Big{|}\ j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\Big{]}$
Since $j^{m}\in J_{\neg Q}$, the agent has no information about $j^{m}$ other than that it is in $J_{\neg Q}$. This means that $j^{m}$ is uniformly distributed across $J_{\neg Q}$, so
$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\right]\geq 1-\frac{1}{|J_{\neg Q}|}\geq 1-\frac{1}{f(T)-g(T)}$
@@
$=$ $\ \Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\cdot\Pr\Big{[}j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
$\geq$ $\ \left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
@@
$\lim_{T\to\infty}\alpha(T)=\lim_{T\to\infty}\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)=(1-0)(1-0-0)=1$
Part 4: Putting it all together. Consider any $\varepsilon\in(0,1]$; to avoid dealing with infinite expectations, we will deal with $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]$ instead of $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$. Since $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq 1$ always, we have
$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]=$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
$+$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})>\varepsilon\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})>\varepsilon\right]$
$\leq$ $\log\varepsilon\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]+1\cdot\left(1-\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]\right)$
$\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
Since $\varepsilon\in(0,1]$, we have $01-\log\varepsilon>$ 0. Also, $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]\geq\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$, so
$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
$\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]$
$\leq$ $\ 1-(1-\log\varepsilon)\alpha(T)$
@@
$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
$=$ $\ \log 1-\operatorname*{\mathbb{E}}_{j^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]$
$\geq$ $\ -1+(1-\log\varepsilon)\alpha(T)$
Therefore
$\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ -1+(1-\log\varepsilon)\lim_{T\to\infty}\alpha(T)$
$\geq$ $\ -1+(1-\log\varepsilon)$
$\geq$ $\ -\log\varepsilon$
This holds for every $\varepsilon\in(0,1]$, which is only possible if $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]=\infty$, as desired. ∎